# Robust Feedback Optimization with Model Uncertainty: A Regularization Approach

### v1

@@ Abstract
Feedback optimization optimizes the steady state of a dynamical system by implementing optimization iterations in closed loop with the plant. It relies on online measurements and limited model information, namely, the input-output sensitivity. In practice, various issues including inaccurate modeling, lack of observation, or changing conditions can lead to sensitivity mismatches, causing closed-loop sub-optimality or even instability. To handle such uncertainties, we pursue robust feedback optimization, where we optimize the closed-loop performance against all possible sensitivities lying in specific uncertainty sets. We provide tractable reformulations for the corresponding min-max problems via regularizations and characterize the online closed-loop performance through the tracking error in case of time-varying optimal solutions. Simulations on a distribution grid illustrate the effectiveness of our robust feedback optimization controller in addressing sensitivity mismatches in a non-stationary environment.
@@ II-A Preliminaries
where $x_{k}\in\mathbb{R}^{n}$, $u_{k}\in\mathbb{R}^{m}$, $y_{k}\in\mathbb{R}^{p}$, $d_{x,k}\in\mathbb{R}^{n}$, and $d_{y,k}\in\mathbb{R}^{p}$ denote the state, input, output, exogenous disturbance, and measurement noise at time $k$, respectively. Further, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, and $C\in\mathbb{R}^{p\times n}$ are system matrices. We focus on a stable system, i.e., the spectral radius $\rho(A)$ of $A$ in (1) is less than $1$. In practice, this condition also holds if this system is prestabilized. Given fixed inputs and disturbances (i.e., $u_{k}=u,d_{x,k}=d_{x},d_{y,k}=d_{y},\forall k\in\mathbb{N}$), system (1) admits a unique steady-state output
@@
To optimize (3), numerical solvers requires an explicit knowledge of the map $y_{\textup{ss}}$ as per (2) with an accurate value of the disturbance $d$, which can be restrictive in applications. In contrast, feedback optimization leverages real-time output measurements and the limited model information, namely, the sensitivity matrix $H$, thus steering system (1) to optimal operating conditions [1, 2].
@@ II-B Example: Detrimental Effects of Inexact Sensitivities
Many practical issues including lack of measurements, corrupted data, and changing conditions cause model uncertainty, i.e., sensitivity errors[2]. We present a motivating example to show how such errors invalidate feedback optimization by inducing closed-loop sub-optimality or instability. While this example is synthetic, we observe similar phenomena in realistic power grid simulations (see Section V).
@@
  • We fix the size of $H$ (i.e., $H\in\mathbb{R}^{3\times 3}$) and add constant perturbations of different intensities. Specifically, $\hat{H}=H+\sigma\Delta_{H}$, where $\sigma\geq 0$, and the elements of $\Delta_{H}$ follow uniform distributions.
@@
To optimize (3), consider the following feedback optimization controller with an inexact $\hat{H}$
$u_{k+1}=u_{k}-2\eta\left(Ru_{k}+\lambda\hat{H}^{\top}Q(y_{k}-r_{k})\right),$ (4)
where $\eta>0$ is the step size. The update (4) follows a gradient descent iteration given the objective (3), using the real-time output measurement $y_{k}$ of (1) to replace the steady-state output $Hu_{k}+d_{k}$. Fig. 1 illustrates the closed-loop performance when the controller (4) is applied to the system (1). We observe from Fig. 1(a) that larger errors in sensitivities cause increased sub-optimality. Furthermore, Fig. 1(b) demonstrates that when $\eta$ is fixed, such a negative effect becomes more pronounced when the problem dimension grows.
@@
Figure 1: Closed-loop performance when the controller (4) with inexact sensitivities is interconnected with the system (2)
@@ II-C Problem Formulation
Motivated by the above observations, we pursue robust feedback optimization, where we optimize a worst-case performance objective given any realization of sensitivity lying in uncertainty sets. In practice, we can obtain through prior knowledge or identification[15, 2] an inexact sensitivity $\hat{H}$, which differs from the true sensitivity $H$ of (1) by $\Delta_{H}$, i.e., $\hat{H}+\Delta_{H}=H$. In view of $\hat{H}$ and the uncertainty $\Delta_{H}$, our robust formulation is
@@
where $\mathcal{D}\subset\mathbb{R}^{p\times m}$ is the uncertainty set wherein $\Delta_{H}$ lies, $d_{k}=C(I\!-\!A)^{-1}d_{x,k}+d_{y,k}$ aggregates the disturbances $d_{x,k}$ and $d_{y,k}$, and $r_{k}$ is the reference at time $k$. Different from (3), in (5) we robustify the steady-state specification of system (1) against the sensitivity uncertainty $\Delta_{H}$. Essentially, (5) implies minimizing the steady-state input-output performance for the worst-case realization of sensitivity. We examine the following types of uncertainty sets.
@@ III-A Tractable Reformulations
We provide tractable reformulations of problem (5), thereby facilitating the design of robust feedback optimization controllers. Let $u_{k}^{*}\in\mathbb{R}^{m}$ be the optimal point of problem (5) at time $k$. We analyze two cases involving the uncertainty sets discussed in Section II-C.
@@
$\rho_{\text{gen}}\!=\!\begin{cases}\varrho_{\text{gen}}\lVert Mu_{k}^{*}\!+\!d_{k}\!-\!r_{k}\rVert/\|u_{k}^{*}\|,\!&\text{if }Mu^{k}_{*}\!+\!d_{k}\!-\!r_{k}\neq 0,\\ \varrho_{\text{gen}}/\|u_{k}^{*}\|,&\text{otherwise}.\end{cases}$ (9)
@@
$\rho_{\textup{col}}=\begin{cases}2\lVert Mu_{k}^{*}+\varepsilon\rVert\varrho_{\textup{col}},&\text{if }Mu_{k}^{*}+\varepsilon\neq 0,\\ \varrho_{\textup{col}},&\text{otherwise}.\end{cases}$ (11)
@@
The objective of (5) can be compactly written as
$\lVert(M+\Delta_{M})u+d_{k}-r_{k}\rVert^{2},$ (12)
where $M=\big{[}R^{\frac{1}{2}};\lambda^{\frac{1}{2}}Q^{\frac{1}{2}}\hat{H}\big{]}\in\mathbb{R}^{(m+p)\times m}$, and $\Delta_{M}=\big{[}0;\lambda^{\frac{1}{2}}Q^{\frac{1}{2}}\Delta_{H}\big{]}\in\mathbb{R}^{(m+p)\times m}$. Since the norm is non-negative, optimizing (12) is equivalent to optimizing $\lVert(M+\Delta_{M})u+d_{k}-r_{k}\rVert$. When $\mathcal{D}=\mathcal{D}_{\text{gen}}$, we perform an analysis similar to that in [32, Theorem 3.1] and obtain
$\max_{\lVert\Delta_{M}\rVert_{F}\leq\varrho_{\text{gen}}}\lVert(M\!+\!\Delta_{M})u\!+\!d_{k}\!-\!r_{k}\rVert=\|Mu+d_{k}-r_{k}\|+\varrho_{\text{gen}}\|u\|.$
@@
$\min_{u\in\mathbb{R}^{m}}~{}$ $\lVert Mu+d_{k}-r_{k}\rVert+\varrho_{\text{gen}}\|u\|,$
$\min_{u\in\mathbb{R}^{m}}~{}$ $\lVert Mu+d_{k}-r_{k}\rVert^{2}+\rho_{\text{gen}}\lVert u\rVert^{2},$
share the same optimal point $u_{k}^{*}$ if the condition (9) holds. This result can be proved by comparing the optimality conditions of both problems and noting that 00 is a subgradient of $\|u\|$ at $u=0$.
@@
$\max_{\begin{subarray}{c}\lVert(\Delta_{M})_{i}\rVert\leq(\varrho_{\textup{col}})_{i}\\ \forall i\in\{1,\dots,m\}\end{subarray}}\lVert(M\!+\!\Delta_{M})u+d_{k}\!-\!r_{k}\rVert\!=\!\lVert Mu+d_{k}\!-\!r_{k}\rVert+\varrho_{\textup{col}}^{\top}|u|.$
@@
$\min_{u\in\mathbb{R}^{m}}~{}\lVert Mu+\varepsilon\rVert+\varrho_{\textup{col}}^{\top}|u|\quad\text{and}\quad\min_{u\in\mathbb{R}^{m}}~{}\lVert Mu+\varepsilon\rVert^{2}+\rho_{\textup{col}}^{\top}|u|$
coincide when the regularizer satisfies the condition (11). ∎
@@ III-B Design of Robust Feedback Optimization Controllers
where $\eta>0$ is the step size. In (14), $\operatorname{\operatorname{prox}}_{\eta\rho_{\textup{col}}}(u)\triangleq\arg\!\min_{u^{\prime}\in\mathbb{R}^{m}}\eta\rho_{\textup{col}}^{\top}|u|+\frac{1}{2}\|u^{\prime}-u\|^{2}$ denotes the proximal operator of $\eta\rho_{\textup{col}}^{\top}|u|$, i.e., soft thresholding $\operatorname{\operatorname{sgn}}(u_{i})\max\{|u_{i}|-\eta(\rho_{\textup{col}})_{i}\}$, where $\operatorname{\operatorname{sgn}}(\cdot)$ is the sign function. Similar to (13), this controller uses the real-time output $y_{k}$ of system (1) and iteratively calculates new inputs.
@@ IV Performance Guarantee
Recall that $u_{k}^{*}$ is the optimal point of problem (5) at time $k$, and that $d_{k}\triangleq C(I\!-\!A)^{-1}d_{x,k}+d_{y,k}$ aggregates the disturbances. We consider stable system (1), i.e., $\rho(A)<1$. Let $x_{\textup{ss},k}\in\mathbb{R}^{n}$ be the steady state of (1) induced by $u_{k}$ and $d_{x,k}$. In other words, $x_{\textup{ss},k}=Ax_{\textup{ss},k}+Bu_{k}+d_{x,k}$, implying $x_{\textup{ss},k}=(I-A)^{-1}(Bu_{k}+d_{x,k})$. For any given positive definite $\bar{Q}\in\mathbb{R}^{n\times n}$, there exists a unique positive definite $P\in\mathbb{R}^{n\times n}$ satisfying the Lyapunov equation $A^{\top}PA-P+\bar{Q}=0$. Let $\|x\|_{P}\triangleq\sqrt{x^{\top}Px}$ be the weighted norm and $\lambda_{\max}(P)$ be the maximum eigenvalue of $P$. Our performance guarantee is as follows.
@@
In Theorem 2, we characterize the closed-loop performance through the joint evolution of the distance $\|u_{k}-u_{k}^{*}\|$ to the optimal point $u_{k}^{*}$ and the distance $\|x_{k}-x_{\textup{ss},k}\|$ to the steady state $x_{\textup{ss},k}$. The upper bound (2) is in the flavor of input-to-state stability [33] and similar to [12, 10, 14, 13]. In contrast to these works, we additionally characterize in (2) the cumulative effects of the given sensitivity uncertainty (i.e., $\|H-\hat{H}\|$) and the regularizer corresponding to the uncertainty set (i.e., $\rho_{\textup{col}}$). The effect of the initial conditions $u_{0}$ and $x_{0}$ vanishes exponentially fast, since $c_{M}\in(0,1)$. The asymptotic error is proportional to the shifts of optimal solutions $u_{k}^{*}$ and disturbances $d_{k}$, as well as the sensitivity uncertainty, i.e., $\|H-\hat{H}\|$. The influence arising from this uncertainty can be tuned via the step size, see the terms in $q_{k}$. It is possible to further establish upper bounds on the distance of the output to the optimal steady-state output through the Lipschitz property of the dynamics (1).
V Numerical Experiments
We present a case study in a distribution grid to showcase the effectiveness of our robust feedback optimization controllers. Specifically, we consider real-time voltage regulation while minimizing active power curtailment and reactive power actuation. Our goal is to show that robustification is effective beyond an academic setting for theoretical guarantees and can address practical challenges such as nonlinear steady states and state-dependent sensitivities.
@@ V Numerical Experiments
Figure 2: Real-time voltage control for a distribution grid after the topology change.

### v2

@@ Abstract
Feedback optimization optimizes the steady state of a dynamical system by implementing optimization iterations in closed loop with the plant. It relies on online measurements and limited model information, namely, the input-output sensitivity. In practice, various issues, including inaccurate modeling, lack of observation, or changing conditions, can lead to sensitivity mismatches, causing closed-loop sub-optimality or even instability. To handle such uncertainties, we pursue robust feedback optimization, where we optimize the closed-loop performance against all possible sensitivities lying in specific uncertainty sets. We provide tractable reformulations for the corresponding min-max problems via regularizations and characterize the online closed-loop performance through the tracking error in case of time-varying optimal solutions. Simulations on a distribution grid illustrate the effectiveness of our robust feedback optimization controller in addressing sensitivity mismatches in a non-stationary environment.
@@ II-A Preliminaries
where $x_{k}\in\mathbb{R}^{n}$, $u_{k}\in\mathbb{R}^{m}$, $y_{k}\in\mathbb{R}^{p}$, $d_{x,k}\in\mathbb{R}^{n}$, and $d_{y,k}\in\mathbb{R}^{p}$ denote the state, input, output, exogenous disturbance, and measurement noise at time $k$, respectively. Further, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, and $C\in\mathbb{R}^{p\times n}$ are system matrices. We focus on a stable system, i.e., the spectral radius $\rho(A)$ of $A$ in (1) lies in $(0,1)$. In practice, this condition also holds if this system is prestabilized by state feedback controllers. Given fixed inputs and disturbances (i.e., $u_{k}=u,d_{x,k}=d_{x},d_{y,k}=d_{y},\forall k\in\mathbb{N}$), system (1) admits a unique steady-state output
@@
To optimize (3), numerical solvers require an explicit knowledge of the map $y_{\textup{ss}}$ as per (2) with an accurate value of the disturbance $d$, which can be restrictive in applications. In contrast, feedback optimization leverages real-time output measurements and the limited model information, namely, the sensitivity matrix $H$, thus steering system (1) to optimal operating conditions [1, 2].
@@ II-B Example: Detrimental Effects of Inexact Sensitivities
Many practical issues including data inadequacy and varying conditions cause model uncertainty, i.e., sensitivity errors[2]. We present a motivating example to show how such errors invalidate feedback optimization by inducing closed-loop sub-optimality or instability. While this example is synthetic, we observe similar phenomena in realistic power grid simulations (see Section V).
@@
  • We fix the size of $H$ (i.e., $H\in\mathbb{R}^{3\times 3}$) and add constant perturbations of different magnitudes. Specifically, $\hat{H}=H+\sigma\Delta_{H}$, where $\sigma\geq 0$, and the elements of $\Delta_{H}$ follow uniform distributions.
@@
To optimize (3) with $r_{k}=0$, consider the following feedback optimization controller using an inexact $\hat{H}$
$u_{k+1}=u_{k}-2\eta\left(Ru_{k}+\lambda\hat{H}^{\top}Qy_{k}\right),$ (4)
where $\eta>0$ is the step size. The update (4) follows a gradient descent iteration given the objective (3), replacing the steady-state output $Hu_{k}+d$ by the real-time output measurement $y_{k}$ of (1). We calculate the optimal value of (3) offline through fmincon in MATLAB with the exact $H$ and $d$. Fig. 1 illustrates the closed-loop optimality gap when the controller (4) is applied to the system (1). We observe from Fig. 1(a) that larger errors in sensitivities cause increased sub-optimality. Furthermore, Fig. 1(b) demonstrates that when $\eta$ is fixed, the detrimental effect becomes more pronounced as the problem dimension grows.
@@
Figure 1: Closed-loop performance when the controller (4) with inexact sensitivities is interconnected with the system (2).
@@ II-C Problem Formulation
Motivated by the above observations, we pursue robust feedback optimization, where we optimize a worst-case performance objective given any realization of sensitivity lying in uncertainty sets. In practice, we can obtain through prior knowledge or identification[15, 2, 31] an inexact sensitivity $\hat{H}$, which differs from the true sensitivity $H$ of (1) by $\Delta_{H}$, i.e., $\hat{H}+\Delta_{H}=H$. In view of $\hat{H}$ and the uncertainty $\Delta_{H}$, our robust formulation is
@@
where $\mathcal{D}\subset\mathbb{R}^{p\times m}$ is the uncertainty set wherein $\Delta_{H}$ lies, $d_{k}=C(I\!-\!A)^{-1}d_{x,k}+d_{y,k}$ aggregates the disturbances $d_{x,k}$ and $d_{y,k}$, and $r_{k}$ is the reference at time $k$. Different from (3), in (5) we robustify the steady-state specification of system (1) against the sensitivity uncertainty $\Delta_{H}$. Essentially, (5) implies minimizing the steady-state input-output performance for the worst-case sensitivity realization. We examine the following types of uncertainty sets.
@@ III-A Tractable Reformulations
We provide tractable reformulations of problem (5), thereby facilitating the design of robust feedback optimization controllers. Let $u_{k}^{*}\in\mathbb{R}^{m}$ be the optimal point of problem (5) at time $k$. The objective function of (5) can be compactly written as
$\Bigg{\lVert}\bigg{(}\underbrace{\begin{bmatrix}R^{\frac{1}{2}}\\ \lambda^{\frac{1}{2}}Q^{\frac{1}{2}}\hat{H}\end{bmatrix}}_{\triangleq M\in\mathbb{R}^{(m\!+\!p)\times m}}+\underbrace{\begin{bmatrix}0\\ \lambda^{\frac{1}{2}}Q^{\frac{1}{2}}\Delta_{H}\end{bmatrix}}_{\triangleq\Delta_{M}\in\mathbb{R}^{(m\!+\!p)\times m}}\bigg{)}u+\underbrace{\begin{bmatrix}0\\ d_{k}-r_{k}\end{bmatrix}}_{\triangleq\varepsilon_{k}\in\mathbb{R}^{m\!+\!p}}\Bigg{\rVert}^{2}.$ (8)
We analyze two cases involving the uncertainty sets discussed in Section II-C.
@@
$\rho_{\text{gen}}\!=\!\begin{cases}\varrho_{\text{gen}}\lVert Mu_{k}^{*}\!+\!\varepsilon_{k}\rVert/\|u_{k}^{*}\|,\!&\text{if }Mu_{k}^{*}\!+\!\varepsilon_{k}\neq 0,\\ \varrho_{\text{gen}}/\|u_{k}^{*}\|,&\text{otherwise},\end{cases}$ (10)
and $M$ and $\varepsilon_{k}$ are given in (8).
@@
$\rho_{\textup{col}}=\begin{cases}2\lVert Mu_{k}^{*}+\varepsilon_{k}\rVert\varrho_{\textup{col}},&\text{if }Mu_{k}^{*}+\varepsilon_{k}\neq 0,\\ \varrho_{\textup{col}},&\text{otherwise}.\end{cases}$ (12)
@@
We analyze the compact representation (8) of the objective of problem (5). In (8), $M=\big{[}R^{\frac{1}{2}};\lambda^{\frac{1}{2}}Q^{\frac{1}{2}}\hat{H}\big{]}\in\mathbb{R}^{(m+p)\times m}$, $\Delta_{M}=\big{[}0;\lambda^{\frac{1}{2}}Q^{\frac{1}{2}}\Delta_{H}\big{]}\in\mathbb{R}^{(m+p)\times m}$, and $\varepsilon_{k}=[0;d_{k}-r_{k}]\in\mathbb{R}^{m+p}$. Since the norm is non-negative, optimizing (8) is equivalent to optimizing $\lVert(M+\Delta_{M})u+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varepsilon_{k}}\rVert$. When $\mathcal{D}=\mathcal{D}_{\text{gen}}$, we perform an analysis similar to that in [33, Theorem 3.1] and obtain
$\max_{\lVert\Delta_{M}\rVert_{F}\leq\varrho_{\text{gen}}}\lVert(M\!+\!\Delta_{M})u\!+\!\varepsilon_{k}\rVert=\|Mu+\varepsilon_{k}\|+\varrho_{\text{gen}}\|u\|.$
@@
$\min_{u\in\mathbb{R}^{m}}\lVert Mu\!+\!{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varepsilon_{k}}\rVert\!+\!\varrho_{\text{gen}}\|u\|~\text{and}~\min_{u\in\mathbb{R}^{m}}\lVert Mu\!+\!{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varepsilon_{k}}\rVert^{2}\!+\!\rho_{\text{gen}}\lVert u\rVert^{2}$
share the same optimal point $u_{k}^{*}$ if the condition (10) holds. This result can be proved by comparing the optimality conditions of both problems and noting that 0 is a subgradient of $\|u\|$ at $u=0$. We further expand the above objective and obtain the equivalent problem (9).
@@
$\max_{\begin{subarray}{c}\lVert(\Delta_{M})_{i}\rVert\leq(\varrho_{\textup{col}})_{i}\\ \forall i\in\{1,\dots,m\}\end{subarray}}\lVert(M\!+\!\Delta_{M})u+\varepsilon_{k}\rVert\!=\!\lVert Mu+\varepsilon_{k}\rVert+\varrho_{\textup{col}}^{\top}|u|.$
@@
$\min_{u\in\mathbb{R}^{m}}~\lVert Mu+\varepsilon_{k}\rVert+\varrho_{\textup{col}}^{\top}|u|\quad\text{and}\quad\min_{u\in\mathbb{R}^{m}}~\lVert Mu+\varepsilon_{k}\rVert^{2}+\rho_{\textup{col}}^{\top}|u|$
coincide when the regularizer satisfies the condition (12). Further expansion of the above objective leads to (11).
@@ III-B Design of Robust Feedback Optimization Controllers
where $\eta>0$ is the step size. In (14), $\operatorname{\operatorname{prox}}_{\eta\rho_{\textup{col}}}(u)\triangleq\arg\!\min_{u^{\prime}\in\mathbb{R}^{m}}\eta\rho_{\textup{col}}^{\top}|u|+\frac{1}{2}\|u^{\prime}-u\|^{2}$ denotes the proximal operator of $\eta\rho_{\textup{col}}^{\top}|u|$, i.e., element-wise soft thresholding $\operatorname{\operatorname{sgn}}(u_{i})\max\{|u_{i}|-\eta(\rho_{\textup{col}})_{i},0\}$, where $\operatorname{\operatorname{sgn}}(\cdot)$ is the sign function. Similar to (13), this controller uses the real-time output $y_{k}$ of system (1) and iteratively calculates new inputs.
@@ IV Performance Guarantee
Recall that $u_{k}^{*}$ is the optimal point of problem (5) at time $k$, and that $d_{k}\triangleq C(I\!-\!A)^{-1}d_{x,k}+d_{y,k}$ aggregates the disturbances. We consider the stable system (1), i.e., $\rho(A)<1$. Let $x_{\textup{ss},k}\in\mathbb{R}^{n}$ be the steady state of (1) induced by $u_{k}$ and $d_{x,k}$. In other words, $x_{\textup{ss},k}=Ax_{\textup{ss},k}+Bu_{k}+d_{x,k}$, implying $x_{\textup{ss},k}=(I-A)^{-1}(Bu_{k}+d_{x,k})$. For any given positive definite $\bar{Q}\in\mathbb{R}^{n\times n}$, there exists a unique positive definite $P\in\mathbb{R}^{n\times n}$ satisfying the Lyapunov equation $A^{\top}PA-P+\bar{Q}=0$. Let $\|x\|_{P}\triangleq\sqrt{x^{\top}Px}$ be the weighted norm and $\lambda_{\max}(P)$ be the maximum eigenvalue of $P$. Our performance guarantee is as follows.
@@
In Theorem 2, we characterize the closed-loop performance through the joint evolution of the distance $\|u_{k}-u_{k}^{*}\|$ to the optimal point $u_{k}^{*}$ and the distance $\|x_{k}-x_{\textup{ss},k}\|$ to the steady state $x_{\textup{ss},k}$. The upper bound (2) is in the flavor of input-to-state stability [34] and similar to [12, 10, 14, 13]. In contrast to these works, we additionally characterize in (2) the cumulative effects of the given sensitivity uncertainty (i.e., $\|H-\hat{H}\|$) and the regularizer corresponding to the uncertainty set (i.e., $\rho_{\textup{col}}$). The effect of the initial conditions $u_{0}$ and $x_{0}$ vanishes exponentially fast, because $c_{M}\in(0,1)$. The asymptotic error is proportional to the shifts of optimal solutions $u_{k}^{*}$ and disturbances $d_{k}$, as well as the sensitivity uncertainty, i.e., $\|H-\hat{H}\|$. The influence arising from this uncertainty can be tuned via the step size, see the terms in $q_{k}$. It is possible to further establish upper bounds on the distance of the output to the optimal steady-state output through the Lipschitz property of the dynamics (1).
@@
Figure 2: This figure illustrates real-time voltage control for a distribution grid after the topology change. The horizontal axis represents the time of day. The curves in the top, middle, and bottom sub-figures indicate the bus voltages, inverter reactive power injections, and inverter active power generations, respectively. The black dashed lines in the top and bottom sub-figures correspond to voltage limits and photovoltaic maximum power points, respectively.
V Numerical Experiments
We present a case study in a distribution grid to showcase the effectiveness of our robust feedback optimization controllers. Specifically, we consider real-time voltage regulation while minimizing active power curtailment and reactive power actuation. Our goal is to show that robustification is effective beyond an academic setting for theoretical guarantees and can address practical challenges such as nonlinear steady states and state-dependent sensitivities. Our code is available at https://github.com/zyhe/robustOFO.